OpenStudy (anonymous):

Admission to the movies costs $8 for adults and $4 for kids. The total bill for your entire extended family was $72. This scenario can be modeled by the equation 8a + 4k = 72. Which is not a possible combination of adults and kids? 3 adults and 12 kids 5 adults and 6 kids 7 adults and 4 kids 8 adults and 2 kids

OpenStudy (anonymous):

Just try out each combination:) 8(3) + 4(12) = 72 8(5) + 4(6) = 64 8(7) + 4(4) = 72 8(8) + 4(2) = 72 Which is not a possible combination of adults and children?

OpenStudy (anonymous):

the 2nd one thanks

OpenStudy (anonymous):

Your welcome:)